OpenStudy (e.cociuba):

Express y as a function of x. The constant C is a positive number. y= In 4x + In C (a) y = 4Cx (b) y = 4x + C (c) y = (4x)^C (d) y = x + 4C Need a walk through plz! :)

OpenStudy (perl):

y is a function of x, you have a typo

OpenStudy (perl):

can you take a screenshot of your worksheet

OpenStudy (perl):

y IS a function of x , dude

OpenStudy (perl):

good luck

OpenStudy (perl):

no, you didnt copy it correctly

OpenStudy (perl):

none of your solutions have ln x in it, notice that?

OpenStudy (perl):

i have a math degree

OpenStudy (perl):

how do you know im not in college? lol

OpenStudy (perl):

you better delete that

OpenStudy (perl):

check your parenthesees or something. the answers dont match the question

OpenStudy (perl):

am i?

OpenStudy (perl):

if y= In 4x + In C y does not equal 4cX

OpenStudy (perl):

well we can get a second opinion if you dont trust me

OpenStudy (perl):

ok bump it

OpenStudy (perl):

you can graph it if you like

OpenStudy (perl):

ln just doesnt disappear

terenzreignz (terenzreignz):

What @perl means is that y is already a function of x. If y appears alone on one side of the equation, with x on the other side, then y is a function of x. :)

OpenStudy (perl):

terenz , his answers dont match the question.

OpenStudy (perl):

e.coiuba lets move on. why live in the past

terenzreignz (terenzreignz):

Oh... right, are you sure there are no logarithms (ln) in your choices, @e.cociuba ?

OpenStudy (perl):

i already posted the solution, but ill do it again -sighs heavily- y= In 4x + In C = ln (4xC) , there we are done

OpenStudy (perl):

i did leave already, but i keep getting notifications

OpenStudy (perl):

im not sure how to unnotify myself

terenzreignz (terenzreignz):

My fault... @perl I'm going to stay out of this, until further notice or instruction...

OpenStudy (perl):

terenz, wait

OpenStudy (perl):

terenz, we need your second opinion

terenzreignz (terenzreignz):

Okay... @e.cociuba You're not going to love this... (probably) but I agree with @perl at almost every point~ y already is expressed as a function of x and ln does not simply disappear... So, if I must have a different opinion... I'm wondering (wildly)... could there have been an expression \(\large e^y\) by any chance?

terenzreignz (terenzreignz):

Well, trust @perl The answers are quite accurate and well-explained :)

OpenStudy (perl):

so we are using log rules here ln (x ) + ln(y) = ln ( x*y) in this case we have ln(4x) + ln (C) = ln ( 4x*C)

terenzreignz (terenzreignz):

None of the answer choices YOU posted are correct. But if there was in fact, a "ln" preceding every choice, then yes, a.) would be correct.

terenzreignz (terenzreignz):

Looking at it right now, methinks :)

OpenStudy (e.cociuba):

Ok nvm boys. Thanks for yalls for ur help :)

terenzreignz (terenzreignz):

So many deleted comments o.O No problem @e.cociuba